OpenStudy (anonymous):

What is the third term of this multiplication pattern? multiplication pattern using the formula 3 • 7n – 1

OpenStudy (anonymous):

3(7n-1) ?

OpenStudy (anonymous):

all you have to do is put n =3 and find the value

OpenStudy (anonymous):

so 3(7(3)-1) = 3(21-1) = 3*20 = 60